OpenStudy (anonymous):

huh? Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (anonymous):

Think about the question. When there are no solutions, what happens with the two linear equations? Do they intersect? Are they same? Or are they parallel to each other. Think about it.

OpenStudy (anonymous):

my head is fried i been working all morning >.<

OpenStudy (anonymous):



OpenStudy (jdoe0001):

Notamathgenius01 ok, can you solve system of 2 equations? by elimination or substitution?

OpenStudy (anonymous):

That's the grid and those are the two equations. When there are no solutions that means the two equations do not intersect. It means they are parallel. when two linear equations are parallel, the slope of those two equations are same!

OpenStudy (anonymous):

im not doing a linear equation??

OpenStudy (jdoe0001):

can you solve system of 2 equations? by elimination or substitution?

OpenStudy (anonymous):

you guys are confusing me

OpenStudy (anonymous):

Well im assuming but, looking at your question, I think the kind of answer should be something like this y=3x+1 and y=3x+7

OpenStudy (jdoe0001):

well, if you don't even know what a system of equations is, I don't see how this exercise applies to you at all

OpenStudy (anonymous):

im Solving Equations with Variables on Both Sides

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

y=3x+1 and y=3x+7 like this??

OpenStudy (anonymous):

Grade 9 math right?

OpenStudy (anonymous):

yes algebra but im NOT working on linear equationss right now >.<

OpenStudy (anonymous):

Hahahaha but equations are actually same as linear equations. So consider my question above again. When you are given two linear equations, how would you find the solution to those two equations?

OpenStudy (anonymous):

i dont know have you seen my username >.<

OpenStudy (anonymous):

im in 11th now i failed algebra and i need help .. im doing a 9th grade class

OpenStudy (anonymous):

i dont know about linear equations

OpenStudy (anonymous):

I will help you through dont worry :)

OpenStudy (anonymous):

So I am going to paraphrase the question. You basically have to create two equations with two variables and there should be no real solutions between those two equations.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no solution is like no answer

OpenStudy (anonymous):

ohh i got ya thanks..

OpenStudy (anonymous):

Yep no solution is lke no answer. So if an equation look like this, y=mx+b and y=nx+e the value of m and n should be same but value of b and e should be different